# Bias-Corrected Data Synthesis for Imbalanced Learning

### v1

@@ 2.2 Multi-Task Imbalanced Learning
In this subsection, we focus on applying bias correction techniques to datasets involving multiple related tasks, a scenario commonly addressed by multi-task learning (MTL). MTL is a machine learning paradigm where multiple tasks are learned simultaneously, enabling the model to leverage shared information and learn a common robust representation (Caruana, 1997; Zhang and Yang, 2021). For example, in genomic studies, researchers analyze gene expression data from different regional populations to identify genetic markers for specific diseases such as Alzheimer’s disease (Zhang and Shen, 2011). In this case, each regional population is regarded as a separate learning task. While the goal of identifying Alzheimer’s disease is shared, the genetic and environmental differences between populations lead to unique data distributions. This makes it necessary to utilize a multi-task learning framework to leverage the common structure. However, the number of individuals with the disease is typically smaller than the number of healthy individuals, creating a within-task imbalance problem (Wu et al., 2018; Guo et al., 2025). We apply the bias correction procedure to such imbalanced MTL problems to improve the predictive performance by leveraging information from all tasks.
@@
where $\sigma(t)=1/(1+e^{-t})$ denotes the logistic function, $\bm{B}\in\mathbb{R}^{d\times r}$ denotes the shared coefficient matrix across $K$ tasks and $\bm{\alpha}_{1},\ldots,\bm{\alpha}_{K}\in\mathbb{R}^{r}$ are task-specific. Denote $\bm{M}=\bm{B}(\bm{\alpha}_{1},\ldots,\bm{\alpha}_{K})\in\mathbb{R}^{d\times K}$ as the unknown coefficient matrix with ${\rm rank}(\bm{M})=r$. Consider the task-specific coefficient vector $\bm{\beta}_{k}=\bm{B}\bm{\alpha}_{k}\in\mathbb{R}^{d}$ as the $k$th column of $\bm{M}$ for $k=1,\ldots,K$. Our goal is to learn the left singular vector space of the shared matrix $\bm{B}$.
@@
For each task, we obtain the estimation $\hat{\bm{\beta}}_{k}^{\rm raw}$ from $(\bm{X}_{ki},Y_{ki})_{i=1}^{n_{k}}$ by minimizing the loss function from the raw data $L^{\rm raw}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta}).$ Consider the minority synthetic samples $(\tilde{\bm{X}}_{ki}^{(1)})_{i=1}^{\tilde{n}_{k1}}$ and majority synthetic samples $(\tilde{\bm{X}}_{ki}^{(0)})_{i=1}^{\tilde{n}_{k0}}$. We can also obtain synthetic and bias-corrected loss functions by
@@
$L^{\rm bc}(\bm{\beta})=$ $\frac{1}{n_{k}+\tilde{n}_{k1}}\left[\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})+\tilde{n}_{k1}\cdot\left\{\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})+\hat{\Delta}_{k0}\right\}\right],$
where $\hat{\Delta}_{k0}$ denotes the bias correction term from the majority samples in the correction set $\mathcal{S}_{kc}$ and majority synthetic samples:
$\hat{\Delta}_{k0}=$ $\frac{1}{|\mathcal{S}_{kc}|}\sum_{i\in\mathcal{S}_{kc}}\ell(\bm{X}_{ki},0;\bm{\beta})-\frac{1}{\tilde{n}_{k0}}\sum_{i=1}^{\tilde{n}_{k0}}\ell(\tilde{\bm{X}}_{ki}^{(0)},0;\bm{\beta}).$
Denote $\hat{\bm{\beta}}_{k}^{\rm syn}$ and $\hat{\bm{\beta}}_{k}^{\rm bc}$ as the minimizers of the synthetic loss $L^{\rm syn}$ and bias-corrected loss $L^{\rm bc}$, respectively.
Next, we consider the estimation of the left singular matrix $\bm{U}$ from any coefficient estimators $(\hat{\bm{\beta}}_{k})_{k=1}^{K}$. First, collect all estimators into $\hat{\bm{M}}=(\hat{\bm{\beta}}_{1},\dots,\hat{\bm{\beta}}_{K})\in\mathbb{R}^{d\times K}$. Next, conduct eigendecomposition of $\hat{\bm{M}}\hat{\bm{M}}^{T}$ such that $\hat{\bm{M}}\hat{\bm{M}}^{T}=\hat{\bm{U}}^{\prime}\hat{\bm{\Lambda}}\hat{\bm{U}}^{\prime\top}$, where $\hat{\bm{U}}^{\prime}\in\mathbb{R}^{d\times d}$ is an orthonormal eigenvector matrix satisfying $(\hat{\bm{U}^{\prime}})^{\top}\hat{\bm{U}}^{\prime}=\bm{I}_{d}$ and $\hat{\bm{\Lambda}}={\rm diag}(\hat{\lambda}_{1},\dots,\hat{\lambda}_{d})$ is a diagonal eigenvalue matrix with decreasing eigenvalues $\hat{\lambda}_{1}\geq\cdots\geq\hat{\lambda}_{d}$. The rank of the latent embedding matrix is estimated by maximizing the eigenvalue ratio such that $\hat{r}=\arg\max_{1\leq r\leq d_{-}}\hat{\lambda}_{r}/\hat{\lambda}_{r+1}$, where $d_{-}<d$ is a constant to avoid the case of extremely small eigenvalues. Finally, take the first $\hat{r}$ columns of $\hat{\bm{U}}^{\prime}$ to obtain the estimated shared embedding matrix $\hat{\bm{U}}=\hat{\bm{U}}^{\prime}_{1:\hat{r}}$. By substituting the estimator $\hat{\bm{\beta}}_{k}$ by the raw estimator $\hat{\bm{\beta}}_{k}^{\rm raw}$, synthetic estimator $\hat{\bm{\beta}}_{k}^{\rm syn}$ and the bias-corrected estimator $\hat{\bm{\beta}}_{k}^{\rm bc}$, we are able to obtain the corresponding latent embedding matrix estimators $\hat{\bm{U}}^{\rm raw}$, $\hat{\bm{U}}^{\rm syn}$ and $\hat{\bm{U}}^{\rm bc}$.
@@
The estimated shared embedding matrix $\hat{\bm{U}}$ helps us to estimate the coefficient in a lower dimension $\hat{r}$ rather than $d$. To obtain the estimation, we can first project the covariates into a lower-dimensional embedding subspace by $\hat{\bm{Z}}_{K+1,i}=\hat{\bm{U}}^{\top}\bm{X}_{K+1,i}$ for each $i=1,\dots,n_{K+1}$. Next, obtain $\hat{\bm{\theta}}_{K+1}$ which minimizes the loss function, for example, $L^{\rm raw}(\bm{\theta})$ on the dataset $(\hat{\bm{Z}}_{K+1,i},Y_{K+1,i})_{i=1}^{n_{K+1}}$. Note that the empirical loss function $L^{\rm raw}$ can be replaced by $L^{\rm syn}$ and $L^{\rm bc}$ depending on the imbalance of task $K+1$. Next, project $\hat{\bm{\theta}}_{K+1}$ back to the coefficient space by $\hat{\bm{\beta}}_{K+1}=\hat{\bm{U}}\hat{\bm{\theta}}_{K+1}$ and obtain the prediction function $\hat{f}_{K+1}(\bm{x})=\sigma(\bm{x}^{\top}\hat{\bm{\beta}}_{K+1})$. By substituting $\hat{\bm{U}}$ with the above $\hat{\bm{U}}^{\rm raw}$, $\hat{\bm{U}}^{\rm syn}$ and $\hat{\bm{U}}^{\rm bc}$, we can obtain the corresponding coefficient estimators and prediction function. In Section 3.2, we provide the theoretical results of the coefficient and embedding matrix estimations of the three methods and show the conditions under which the bias correction procedure outperforms the synthetic procedure.
@@ 2.3 Average Treatment Effect Estimation
Suppose $Y(1)$ and $Y(0)$ are the potential responses under treatment $Z=1$ and control $Z=0$. The observed response is a function of the potential responses and the treatment indicator:
$Y=ZY(1)+(1-Z)Y(0).$
Then the ATE is defined as
$\tau=\mathbb{E}\{Y(1)\}-\mathbb{E}\{Y(0)\}.$
@@
$\hat{\mu}_{1}^{\rm AIPW}=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\frac{Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}}{e(\bm{X}_{i})}+\mu_{1}(\bm{X}_{i})\right],$ (2.11)
$\hat{\mu}_{0}^{\rm AIPW}=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\frac{(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}}{1-e(\bm{X}_{i})}+\mu_{0}(\bm{X}_{i})\right],$ (2.12)
$\hat{\tau}^{\rm AIPW}=$ $\hat{\mu}_{1}^{\rm AIPW}-\hat{\mu}_{0}^{\rm AIPW}.$ (2.13)
With the observations $(\bm{X}_{i},Y_{i},Z_{i})_{i=1}^{n}$, we can first fit separate regression models of the responses on the covariates and obtain the coefficient estimators $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{0}$ under treatment and control, respectively. Let $\hat{\mu}_{1}(\bm{x})=\bm{x}^{\top}\hat{\bm{\beta}}_{1}$ and $\hat{\mu}_{0}(\bm{x})=\bm{x}^{\top}\hat{\bm{\beta}}_{0}$ be the estimated responses under treatment and control, respectively.
Next, we consider the propensity score estimation. Suppose we are interested in a loss function $\ell_{f}$, e.g., logistic loss, for a prediction function $f:\mathbb{R}^{d}\to[0,1]$. With the raw data, we can obtain the propensity score estimation $\hat{f}^{\rm raw}$ by minimizing the empirical loss function $L^{\rm raw}$ in (2.1). Assume that there are $\tilde{n}_{1}$ synthetic minority covariate samples $\tilde{\bm{X}}_{i}^{(1)}$, then we can obtain the synthetic-augmented propensity score estimation $\hat{f}^{\rm syn}$ by minimizing the $L^{\rm syn}$ in (2.2). Partition the control index set into a generation set $\mathcal{S}_{0g}$ and a correction set $\mathcal{S}_{0c}$. Applying the same synthetic generator to obtain $\tilde{n}_{0}$ control synthetic covariates $\tilde{\bm{X}}_{i}^{(0)}$ from the generation set, we can obtain the bias-corrected propensity score estimation $\hat{f}^{\rm bc}$ by minimizing $L^{\rm bc}$ in (2.8).
Plugging the treatment and control models $\hat{\mu}_{1}(\bm{x})$ and $\hat{\mu}_{0}(\bm{x})$ as well as the propensity score estimation $\hat{e}(\bm{x})\in\{\hat{f}^{\rm raw}(\bm{x}),\hat{f}^{\rm syn}(\bm{x}),\hat{f}^{\rm bc}(\bm{x})\}$ into (2.13), the AIPW estimators $\hat{\tau}^{\rm raw},\hat{\tau}^{\rm syn}$ and $\hat{\tau}^{\rm bc}$ can be derived corresponding to the propensity score estimators.
@@ 3.1 Bias Correction for Risk Functions
In this subsection, we first propose an upper bound for the difference between the minority and majority bias correction terms, $|\hat{\Delta}_{1}-\hat{\Delta}_{0}|$. This result provides a theoretical guarantee for the construction of the bias-corrected loss function $L^{\rm bc}$ in (2.8). We then derive a lower bound for the minority bias correction term $\hat{\Delta}_{1}$ of SMOTE, thereby illustrating that treating synthetic samples as real data may introduce bias. Finally, we present the properties of the bias-corrected predictor with respect to the balanced population risk function, demonstrating that the proposed procedure effectively leverages synthetic samples while ensuring strong performance on imbalanced data.
Assumption 1.
Consider the support of covariates $\mathcal{X}\subseteq\mathbb{R}^{d}$. Suppose the following conditions are satisfied:
@@
    $\mathcal{P}_{1}=(\mathcal{P}_{0})_{\#T},\quad\mathcal{W}_{1}(\tilde{\mathcal{P}}_{1},(\tilde{\mathcal{P}}_{0})_{\#T})\leq\varepsilon_{T},$
    where $(\cdot)_{\#T}$ denotes the distribution transformation pushforward by $T$. For instance, if $\bm{X}\sim\mathcal{P}_{0}$, then $T(\bm{X})\sim(\mathcal{P}_{0})_{\#T}$.
@@
    Lipschitz smoothness of the transformed loss. There exist constants $L_{\ell},L_{T}>0$ such that for all $\bm{x}_{1},\bm{x}_{2}\in\mathcal{X}$ and $y\in\{0,1\}$,
    $|\ell_{f}(T(\bm{x}_{1}),y)-\ell_{f}(T(\bm{x}_{2}),y)|\leq$ $L_{T}\|\bm{x}_{1}-\bm{x}_{2}\|_{2},$

### v2

@@ 2.2 Multi-Task Imbalanced Learning
In this subsection, we focus on applying bias correction techniques to datasets involving multiple related tasks, a scenario commonly addressed by multi-task learning (MTL). MTL is a machine learning paradigm where multiple related tasks are learned simultaneously, enabling the model to leverage shared information and learn a common robust representation (Caruana, 1997; Zhang and Yang, 2021). For example, in genomic studies, researchers analyze gene expression data from different regional populations to identify genetic markers for specific diseases such as Alzheimer’s disease (Zhang and Shen, 2011). In this case, each regional population is regarded as a separate learning task. While the goal of identifying Alzheimer’s disease is shared, the genetic and environmental differences between populations lead to unique data distributions. This makes it necessary to utilize a multi-task learning framework to leverage the common structure. However, the number of individuals with the disease is typically smaller than the number of healthy individuals, creating a within-task imbalance problem (Wu et al., 2018; Guo et al., 2025). We apply the bias correction procedure to such imbalanced MTL problems to improve the predictive performance by leveraging information from all tasks.
@@
where $\sigma(t)=1/(1+\exp(-t))$ denotes the logistic function, $\bm{B}\in\mathbb{R}^{d\times r}$ denotes the shared coefficient matrix across $K$ tasks and $\bm{\alpha}_{1},\ldots,\bm{\alpha}_{K}\in\mathbb{R}^{r}$ are task-specific latent loading vectors. Denote $\bm{M}=\bm{B}(\bm{\alpha}_{1},\ldots,\bm{\alpha}_{K})\in\mathbb{R}^{d\times K}$ as the unknown coefficient matrix with ${\rm rank}(\bm{M})=r$. Consider the task-specific coefficient vector $\bm{\beta}_{k}=\bm{B}\bm{\alpha}_{k}\in\mathbb{R}^{d}$ as the $k$th column of $\bm{M}$ for $k=1,\ldots,K$. Our goal is to learn the left singular vector space of the shared matrix $\bm{B}$.
@@
For each task, we obtain the estimation $\widehat{\bm{\beta}}_{k}^{\rm raw}$ from $(\bm{X}_{ki},Y_{ki})_{i=1}^{n_{k}}$ by minimizing the loss function from the raw data $L^{\rm raw}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta}).$ Consider the minority synthetic samples $(\tilde{\bm{X}}_{ki}^{(1)})_{i=1}^{\tilde{n}_{k1}}$ and majority synthetic samples $(\tilde{\bm{X}}_{ki}^{(0)})_{i=1}^{\tilde{n}_{k0}}$. We can also obtain synthetic and bias-corrected loss functions by
@@
$L^{\rm bc}(\bm{\beta})=$ $\frac{1}{n_{k}+\tilde{n}_{k1}}\left[\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})+\tilde{n}_{k1}\cdot\left\{\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})+\widehat{\Delta}_{k0}\right\}\right],$
where $\widehat{\Delta}_{k0}$ denotes the bias correction term from the majority samples in the correction set $\mathcal{S}_{kc}$ and majority synthetic samples:
$\widehat{\Delta}_{k0}=$ $\frac{1}{|\mathcal{S}_{kc}|}\sum_{i\in\mathcal{S}_{kc}}\ell(\bm{X}_{ki},0;\bm{\beta})-\frac{1}{\tilde{n}_{k0}}\sum_{i=1}^{\tilde{n}_{k0}}\ell(\tilde{\bm{X}}_{ki}^{(0)},0;\bm{\beta}).$
Denote $\widehat{\bm{\beta}}_{k}^{\rm syn}$ and $\widehat{\bm{\beta}}_{k}^{\rm bc}$ as the minimizers of the synthetic loss $L^{\rm syn}$ and bias-corrected loss $L^{\rm bc}$, respectively.
Next, we consider the estimation of the left singular matrix $\bm{U}$ from any coefficient estimators $(\widehat{\bm{\beta}}_{k})_{k=1}^{K}$. First, collect all estimators into $\widehat{\bm{M}}=(\widehat{\bm{\beta}}_{1},\dots,\widehat{\bm{\beta}}_{K})\in\mathbb{R}^{d\times K}$. Next, conduct eigendecomposition of $\widehat{\bm{M}}\widehat{\bm{M}}^{T}$ such that $\widehat{\bm{M}}\widehat{\bm{M}}^{T}=\widehat{\bm{U}}^{\prime}\widehat{\bm{\Lambda}}\widehat{\bm{U}}^{\prime\top}$, where $\widehat{\bm{U}}^{\prime}\in\mathbb{R}^{d\times d}$ is an orthonormal eigenvector matrix satisfying $(\widehat{\bm{U}^{\prime}})^{\top}\widehat{\bm{U}}^{\prime}=\bm{I}_{d}$ and $\widehat{\bm{\Lambda}}={\rm diag}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{d})$ is a diagonal eigenvalue matrix with decreasing eigenvalues $\widehat{\lambda}_{1}\geq\cdots\geq\widehat{\lambda}_{d}$. The rank of the latent embedding matrix is estimated using an eigenvalue ratio criterion that has been widely adopted for determining the number of latent factors in high-dimensional spectral and factor models (Ahn and Horenstein, 2013; Fan et al., 2019; Wang et al., 2026). Specifically, the rank is selected by maximizing the eigenvalue ratio such that $\widehat{r}=\arg\max_{1\leq r\leq d_{-}}\widehat{\lambda}_{r}/\widehat{\lambda}_{r+1}$, where $d_{-}<d$ is a constant to avoid numerical instability caused by extremely small eigenvalues. Finally, take the first $\widehat{r}$ columns of $\widehat{\bm{U}}^{\prime}$ to obtain the estimated shared embedding matrix $\widehat{\bm{U}}=\widehat{\bm{U}}^{\prime}_{1:\widehat{r}}$. Applying the same construction to different coefficient estimators yields the corresponding latent embedding matrix estimators. In particular, using the raw estimator $\widehat{\bm{\beta}}_{k}^{\rm raw}$, the synthetic estimator $\widehat{\bm{\beta}}_{k}^{\rm syn}$, and the bias-corrected estimator $\widehat{\bm{\beta}}_{k}^{\rm bc}$ in place of $\widehat{\bm{\beta}}_{k}$ results in $\widehat{\bm{U}}^{\rm raw}$, $\widehat{\bm{U}}^{\rm syn}$, and $\widehat{\bm{U}}^{\rm bc}$, respectively.
@@
The estimated shared embedding matrix $\widehat{\bm{U}}$ helps us to estimate the coefficient in a lower dimension $\widehat{r}$ rather than $d$. To obtain the estimation, we can first project the covariates into a lower-dimensional embedding subspace by $\widehat{\bm{Z}}_{K+1,i}=\widehat{\bm{U}}^{\top}\bm{X}_{K+1,i}$ for each $i=1,\dots,n_{K+1}$. Next, obtain $\widehat{\bm{\theta}}_{K+1}$ which minimizes the loss function, for example, $L^{\rm raw}(\bm{\theta})$ on the dataset $(\widehat{\bm{Z}}_{K+1,i},Y_{K+1,i})_{i=1}^{n_{K+1}}$. Note that the empirical loss function $L^{\rm raw}$ can be replaced by $L^{\rm syn}$ and $L^{\rm bc}$ depending on the imbalance of task $K+1$. Then project $\widehat{\bm{\theta}}_{K+1}$ back to the coefficient space by $\widehat{\bm{\beta}}_{K+1}=\widehat{\bm{U}}\widehat{\bm{\theta}}_{K+1}$ and obtain the prediction function $\widehat{f}_{K+1}(\bm{x})=\sigma(\bm{x}^{\top}\widehat{\bm{\beta}}_{K+1})$. By substituting $\widehat{\bm{U}}$ with the above $\widehat{\bm{U}}^{\rm raw}$, $\widehat{\bm{U}}^{\rm syn}$ and $\widehat{\bm{U}}^{\rm bc}$, we can obtain the corresponding coefficient estimators and prediction function. In Section 3.2, we provide the theoretical results of the coefficient and embedding matrix estimations of the three methods and show the conditions under which the bias correction procedure outperforms the synthetic procedure.
@@ 2.3 Average Treatment Effect Estimation
Suppose $Y(1)$ and $Y(0)$ are the potential responses under treatment $Z=1$ and control $Z=0$. The observed response is a function of the potential responses and the treatment indicator: Y=ZY(1)+(1Z)Y(0)Y=ZY(1)+(1-Z)Y(0). Then the ATE is defined as $\tau=\mathbb{E}\{Y(1)\}-\mathbb{E}\{Y(0)\}$.
@@
$\widehat{\mu}_{1}^{\rm AIPW}=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\frac{Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}}{e(\bm{X}_{i})}+\mu_{1}(\bm{X}_{i})\right],$ (2.11)
$\widehat{\mu}_{0}^{\rm AIPW}=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\frac{(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}}{1-e(\bm{X}_{i})}+\mu_{0}(\bm{X}_{i})\right],$ (2.12)
$\widehat{\tau}^{\rm AIPW}=$ $\widehat{\mu}_{1}^{\rm AIPW}-\widehat{\mu}_{0}^{\rm AIPW}.$ (2.13)
With the observations $(\bm{X}_{i},Y_{i},Z_{i})_{i=1}^{n}$, we can first fit separate regression models of the responses on the covariates and obtain the coefficient estimators $\widehat{\bm{\beta}}_{1}$ and $\widehat{\bm{\beta}}_{0}$ under treatment and control, respectively. Let $\widehat{\mu}_{1}(\bm{x})=\bm{x}^{\top}\widehat{\bm{\beta}}_{1}$ and $\widehat{\mu}_{0}(\bm{x})=\bm{x}^{\top}\widehat{\bm{\beta}}_{0}$ be the estimated responses under treatment and control, respectively.
Next, we consider the propensity score estimation. Suppose we are interested in a loss function $\ell_{f}$, e.g., logistic loss, for a prediction function $f:\mathbb{R}^{d}\to(0,1)$. With the raw data, we can obtain the propensity score estimation $\widehat{f}^{\rm raw}$ by minimizing the empirical loss function $L^{\rm raw}$ in (2.1). Assume that there are $\tilde{n}_{1}$ synthetic minority covariate samples $\tilde{\bm{X}}_{i}^{(1)}$, then we can obtain the synthetic-augmented propensity score estimation $\widehat{f}^{\rm syn}$ by minimizing the $L^{\rm syn}$ in (2.2). Partition the control index set into a generation set $\mathcal{S}_{0g}$ and a correction set $\mathcal{S}_{0c}$. Applying the same synthetic generator to obtain $\tilde{n}_{0}$ control synthetic covariates $\tilde{\bm{X}}_{i}^{(0)}$ from the generation set, we can obtain the bias-corrected propensity score estimation $\widehat{f}^{\rm bc}$ by minimizing $L^{\rm bc}$ in (2.8).
Plugging the treatment and control models $\widehat{\mu}_{1}(\bm{x})$ and $\widehat{\mu}_{0}(\bm{x})$ as well as the propensity score estimation $\widehat{e}(\bm{x})\in\{\widehat{f}^{\rm raw}(\bm{x}),\widehat{f}^{\rm syn}(\bm{x}),\widehat{f}^{\rm bc}(\bm{x})\}$ into (2.13), the AIPW estimators $\widehat{\tau}^{\rm raw},\widehat{\tau}^{\rm syn}$ and $\widehat{\tau}^{\rm bc}$ can be derived corresponding to the propensity score estimators.
@@ 3.1 Bias Correction for Risk Functions
In this subsection, we first propose an upper bound for the difference between the minority and majority bias correction terms, $|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|$. This result provides a theoretical guarantee for the construction of the bias-corrected loss function $L^{\rm bc}$ in (2.8). We then derive a lower bound for the minority bias correction term $\widehat{\Delta}_{1}$ of SMOTE, thereby illustrating that treating synthetic samples as real data may introduce bias. Finally, we present the properties of the bias-corrected predictor with respect to the balanced population risk function, demonstrating that the proposed procedure effectively leverages synthetic samples while ensuring strong performance on imbalanced data.
To establish finite-sample guarantees for the proposed bias-corrected loss and predictor, we impose a set of standard regularity conditions. These assumptions are introduced to control the behavior of the loss function, the complexity of the classifier class, and the dependence structure of synthetic samples. They are commonly adopted in empirical risk minimization and learning with synthetic or augmented data, and do not impose algorithm-specific restrictions.
We first impose a boundedness condition on the prediction function and the associated loss. This assumption ensures that the empirical risk and its bias correction terms are well-defined and uniformly controlled, which is essential for deriving concentration inequalities and uniform deviation bounds. Such boundedness conditions are standard in the analysis of classification methods with probabilistic outputs.
Assumption 3.1 (Bounded prediction and loss).
There exists a constant $\varepsilon\in(0,1/2)$ such that for all $f\in\mathcal{F}$ and all $\bm{x}\in\mathcal{X}$, $f(\bm{x})\in[\varepsilon,1-\varepsilon]$ and there exists a constant $M_{\varepsilon}>0$ such that $0\leq\ell_{f}(\bm{x},y)\leq M_{\varepsilon}$ for all $(\bm{x},y)$. For instance, for the binary cross-entropy loss $\ell_{f}(\bm{x},y)=-y\log f(\bm{x})-(1-y)\log(1-f(\bm{x}))$, we have $M_{\varepsilon}=-\log\varepsilon$.
Next, we control the richness of the classifier class through its induced loss class. Bounding the Rademacher complexity of the normalized loss functions allows us to quantify the uniform generalization error of empirical risk minimization over $\mathcal{F}$. This assumption provides a standard measure of function class complexity and plays a central role in our finite-sample analysis.
Assumption 3.2 (Function class complexity).
Let $\mathcal{L}=\{(\bm{x},y)\mapsto\ell_{f}(\bm{x},y)/M_{\varepsilon}:f\in\mathcal{F}\}$. The Rademacher complexity of a function class $\mathcal{L}$ with sample size $m\geq 1$ is defined as
$\mathfrak{R}_{m}(\mathcal{L}):=\mathbb{E}_{\bm{Z}_{1:m},\,\bm{\sigma}}\left[\sup_{g\in\mathcal{L}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\,g(\bm{Z}_{i})\right],$
where $\bm{Z}_{1},\ldots,\bm{Z}_{m}$ are i.i.d. observations with $\bm{Z}_{i}=(\bm{X}_{i},Y_{i})$, $\sigma_{1},\ldots,\sigma_{m}$ are i.i.d. Rademacher random variables satisfying $\mathbb{P}(\sigma_{i}=1)=\mathbb{P}(\sigma_{i}=-1)=1/2$, and $(\sigma_{i})_{i=1}^{m}$ are independent of $(\bm{Z}_{i})_{i=1}^{m}$. There exists a constant $C_{\mathcal{F}}>0$ such that for all $m\geq 1$, the Rademacher complexity satisfies $\mathfrak{R}_{m}(\mathcal{L})\leq C_{\mathcal{F}}/\sqrt{m}$.
We now formalize the stochastic structure of the synthetic samples. Rather than assuming unconditional independence, we introduce a conditional independence framework in which synthetic data are generated from a random generator $\mathcal{G}$ and are independent of the evaluation samples given $\mathcal{G}$. This formulation accommodates many synthetic data generation mechanisms, including resampling and model-based procedures, while preserving tractable theoretical analysis.
Assumption 3.3 (Synthetic conditional independence).
There exists a random object $\mathcal{G}$ (the synthetic data generator) such that conditional on $\mathcal{G}$, the synthetic samples $\{\tilde{\bm{X}}_{i}^{(1)}\}_{i=1}^{\tilde{n}_{1}}$ are i.i.d. from $\tilde{\mathcal{P}}_{1}$ and $\{\tilde{\bm{X}}_{i}^{(0)}\}_{i=1}^{\tilde{n}_{0}}$ are i.i.d. from $\tilde{\mathcal{P}}_{0}$. Moreover, conditional on $\mathcal{G}$, the synthetic samples are independent of the observed samples used to evaluate $L^{\rm bc}$ and $\widehat{\Delta}_{0}$.
Finally, to characterize the structural relationship between the minority and majority populations, we introduce an alignment assumption based on distributional transformations. This assumption is imposed at an abstract level to quantify the degree of distributional similarity between populations, rather than to require an explicit or computable transformation in practice. It serves as an analytical device for controlling the discrepancy between bias correction terms across populations and for formalizing the effect of synthetic samples under distributional shift.
Assumption 3.4.
Consider the support of covariates $\mathcal{X}\subseteq\mathbb{R}^{d}$. Suppose the following conditions are satisfied uniformly for all $f\in\mathcal{F}$:
@@
    $\mathcal{W}_{1}(\mathcal{P}_{1},(\mathcal{P}_{0})_{\#T})\leq\varepsilon_{T},\quad\text{and }\quad\mathcal{W}_{1}(\tilde{\mathcal{P}}_{1},(\tilde{\mathcal{P}}_{0})_{\#T})\leq\varepsilon_{T}\quad\text{a.s. conditional on }\mathcal{G},$
    where $(\cdot)_{\#T}$ denotes the pushforward distribution induced by transformation $T$, and $\mathcal{W}_{1}(\cdot,\cdot)$ represents the Wasserstein–1 distance.
@@
    Lipschitz smoothness of the transformed loss. There exists constant $L_{\ell}>0$ such that for all $\bm{x}_{1},\bm{x}_{2}\in\mathcal{X}$ and $y\in\{0,1\}$,
    $|\ell_{f}(T(\bm{x}_{1}),y)-\ell_{f}(T(\bm{x}_{2}),y)|\leq$ $L_{\ell}\|\bm{x}_{1}-\bm{x}_{2}\|_{2},$